OpenStudy (anonymous):

Use quotient rule to simplify: square root of (7/100)

OpenStudy (anonymous):

√7/10

OpenStudy (anonymous):

\[\sqrt{\frac{7}{10}}=\frac{\sqrt{7}}{\sqrt{10}}\]

OpenStudy (anonymous):

How about sqrt of 2/25?

OpenStudy (anonymous):

100's not 10

OpenStudy (anonymous):

√2/5

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{\sqrt{25}}=\frac{\sqrt{2}}{5}\]

OpenStudy (anonymous):

brilliant, thakns you!

myininaya (myininaya):

\[\sqrt{\frac{7}{100}}=\frac{\sqrt{7}}{\sqrt{100}}=\frac{\sqrt{7}}{10}\]

OpenStudy (anonymous):

yes thats what i was going for^

myininaya (myininaya):

lol :)